USE THE PROPERTIES OFLOGARITHMS TO WRITE THE FOLLOWING AS A SUM, DIFFERENCE,AND OR MULTIPLE OF LOGARITHMS:
LOG78x
log5 y/x
log 4 x^5

Answer by stanbon(48535)

USE THE PROPERTIES OF LOGARITHMS TO WRITE THE FOLLOWING AS A SUM, DIFFERENCE,AND OR MULTIPLE OF LOGARITHMS:
-----------------
LOG78x
Is 7 the base or part of the inverse-log ?
----------------------------------------------
log5 y/x = log5(y) - log5(x)
----------------------------------------------
log 4 x^5 = 5log4 (x)
================================================
Cheers,
Stan H.